OpenStudy (anonymous):

let f(x)=24/3x-2 find f(-2)

OpenStudy (anonymous):

f(-2) is the value of the function when x = -2. So what's the value of the function, 24/3x - 2, when x = -2?

OpenStudy (anonymous):

-6 3 -3 6

OpenStudy (anonymous):

What is your answer?

OpenStudy (anonymous):

i think it -6

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

well, what is in your denominator, 3x or (3x - 2)?

OpenStudy (anonymous):

3x-2

OpenStudy (anonymous):

That is, is the function 24/(3x) - 2 or 24/(3x-2)?

OpenStudy (anonymous):

so if x = -2, 3x - 2 = 3(-2) - 2 = -6 - 2 = -8

OpenStudy (anonymous):

ok so now i divide 24/-8=-3

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

easyaspi314 broke it down correctly

OpenStudy (anonymous):

im sorry, -3